# Secure Targeted Message Dissemination in IoT Using Blockchain Enabled Edge Computing

### v1

@@ Abstract
Smart devices are considered as an integral part of Internet of Things (IoT), have an aim to make a dynamic network to exchange information, collect data, analysis, and make optimal decisions in an autonomous way to achieve more efficient, automatic, and economical services. Message dissemination among these smart devices allows adding new features, sending updated instructions, alerts or safety messages, informing the pricing information or billing amount, incentives, and installing security patches. On one hand, such message disseminations are directly beneficial to the all parties involved in the IoT system. On the other hand, due to remote procedure, smart devices, vendors, and other involved authorities might have to meet a number of security, privacy, and performance related concerns while disseminating messages among targeted devices. To this end, in this paper, we design STarEdgeChain, a security and privacy aware targeted message dissemination in IoT to show how blockchain along with advanced cryptographic techniques are devoted to address such concerns. In fact, the STarEdgeChain employs a permissioned blockchain assisted edge computing in order to expedite a single signcrypted message dissemination among targeted groups of devices, at the same time avoiding the dependency of utilizing multiple unicasting approaches. Finally, we develop a software prototype of STarEdgeChain and show it’s practicability for smart devices. The codes are publicly available at https://github.com/mbaqer/Blockchain-IoT
@@ I Introduction
The Internet of Things (IoT) [1, 2, 3], has been rapidly growing more than ever before. From consumer electronic devices to industry, the IoT is making it possible to connect a variety of things and people to the Internet, and each other to bring enhancements to our quality of lives, surroundings, and system performances. Such things, also called smart devices (SDs) connect to the Internet and each other to exchange information, collect data, analyze, and make better decisions in an autonomous way without human operation. These abilities make SDs very popular in various application areas in the home, smart power grid, building, city, industry, healthcare, and many more. In specific, personal and home devices connected with IoT [4] have been receiving much attention in the consumer electronics landscape since it is expected that the IoT is one of the key enablers to bring simplicity and ubiquitous connectivity to the consumers.
@@
Paper Organization: The remainder of this paper can be summarized as follows. Section II describes the problem statement including system settings, adversary model, and security requirements. The following section, the preliminary background of our construction is considered. Next, section IV presents the detailed construction of the proposed STarEdgeChain. After that, section V summarizes the security and performance analysis. Section VI discusses the related works. Finally, VII concludes the presented work. The list of notations and their definitions used in this paper is described in Table I.
@@
$\lambda$
Security parameter
$p$
Prime number
$\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{3}$
Three cyclic groups
$e$
Bilinear map
$g_{1},g_{2}$
Generator of $\mathbb{G}_{1}$, generator of $\mathbb{G}_{2}$
$\mathbb{Z}_{p}$
Integers modulo $p$
$pk$
Public key
$sk$
Asymmetric secret key
$mk$
Master key
$key_{sym}$
Symmetric secret key
$key_{sign}$
Signing key
$key_{ver}$
Verification key
$msg$
Plaintext message
$\mathcal{S}$
The set of attributes
$\mathcal{T}$
Access tree structure
$\mathcal{A}$
An adversary
$\mathcal{ST}$
Signed Ciphertext
$\mathcal{CT}_{msg}$
Ciphertext of message
$\{0,1\}^{*}$
Set of binary strings
$\mathcal{Y}$
Set of leaf nodes in access tree
$\mathcal{H}_{1},\mathcal{H}_{2}$
Hash functions
$\pi$
Signature
Concatenation operator
$\oplus$
XOR operation
@@ III-D Permissioned Blockchain and Consensus
Blockchain [27, 28, 29, 30, 31], a tamper-proof and consensus based distributed ledger technique, consists of a chain of blocks. The security and privacy of blockchain is ensured by cryptographic techniques, consensus, and incentive mechanisms. Blockchain offers data auditability by employing authenticated blocks attached in the growing chain. In the chain, each block keeps records of financial transactions or non-financial data, and a number of blocks form a chain by including the cryptographic hash of the previous block.
@@ IV-D Message Dissemination
A discussion on the proposed permissioned blockchain architecture is as follows in terms of the following three point of views.
@@
(iii) Blockchain Storage in Lightweight Nodes: Usually, the lightweight IoT smart device nodes have very limited storage to keep the entire blockchain and update it over time. Therefore, the interesting point is that in our permissioned blockchain, the lightweight nodes do not required to store the blockchain, rather the nodes will be relied on edge computing nodes.
@@ V-B Implementation and Performance Evaluation
Implementation Environment: The service providers, PoA validators, and edge computing nodes commonly have high computational devices relatively to IoT smart devices. Thus, as presented in Fig. 3, we have taken two laptop computer, representing as two host machines (Acer Predator Triton 500 and Lenovo IdeaPad 320S) as service providers, PoA validators, and edge computing nodes, whereas, one Raspberry Pi device (RPi 3 Model B+) as IoT smart device. In the host machines, we have configured Oracle virtual box (https://www.virtualbox.org/) and installed Ubuntu 18.04 with one core, RAM 4GB as virtual machines. In fact, we have considered the aforementioned virtual machines and one smart device as node 1, 2, and 3, respectively. Besides, we have used Python 3.6 version for developing the software prototype. After that, we have used the bridge network feature from Oracle virtual box and made a local network among the host machines, virtual machines, and the RPi device for the purpose of on-chain implementation. We have also developed a graphical user interface (GUI) for convenience operation. Specifically, in Fig. 3, the node 1 represents a service provider making transactions and doing PoA self-validation of transactions to add in the chain. On the other hand, node 2 shows an edge computing node connected with a blockchain network and supported to perform targeted message dissemination to the RPi device considered as IoT smart device in node 3.
@@
Node 1
Intel Core i7-10th generation, RAM 32GB, and Microsoft Windows 10 operating system)
Node 2
Intel Core i5-8th generation, RAM 8GB, and Microsoft Windows 10 operating system)
Node 3
ARM Cortex-A53 (ARMv8) 1.4GHz quad-core processor, 1GB RAM, and Raspberry Pi OS)
Libraries
Flask, Tkinter, Charm-Crypto, hashlib
IP and Port addresses
Node 1 VM1: 192.168.100.10:5000,
@@ VI Related Works
Typical Message Dissemination: In recent years, many solutions on typical message dissemination approaches have been introduced. For instance, the work in [36] focuses on secure and privacy-aware emergency vehicular message dissemination named as BTMPP. This solution utilizes bloom filter enabled private set intersection techniques to offer strong conditional privacy. Also, this proposed approach also supports precise trust management simultaneously. Moreover, the authors in [37] proposes a privacy preserving trust management approach for message dissemination in vehicular space-air-ground integrated networks, which has an aim to ensure strong applicability and robustness by employing 0-1 encoding approach. A novel low-latency cooperative security method presents in [38] for platoon-assisted vehicular network to ensure timely and secure vehicular data dissemination. Here, a secret key agreement scheme is employed, in which the secret keys are generated based on the correlation of vehicles radio channels. The work in [39] introduces an end-to-end fast and secure firmware update framework IoT device like embedded systems. This work mainly employed a hardware-embedded cryptographic primitive referred to as a physical unclonable function (PUF) technique. Another work presented in [40] proposes a remote attestation-based firmware update for IoT. Kornaros et al. [41] presents a secure firmware update for connected vehicles by including timestamp, version of update as well signed metadata of update.
@@
Integration of Blockchain: In terms of blockchain integration in message dissemination, existing works mainly employ blockchain along with different cryptographic schemes. In this context, cryptographic techniques are utilized to guarantee the security and privacy related requirements to the involved entities and message contents. For example, the work in [47] constructs a public key infrastructure (PKI) based private blockchain to offer privacy-aware message dissemination in vehicular networks. Another work in [48], the authors develop a semi-decentralized attribute-based blockchain to distribute the traffic signal messages and assist traffic control among the vehicles with high availability and privacy preserving ways in IoV. Here, the authors utilize blockchain along with a ciphertext-policy ABE scheme. Besides, the authors in [49] and [50] presents on how typical blockchain structure is able to handle data dissemination in Industrial IoT and Internet of Drones, respectively. Moreover, Baza et al. [51] leverages attribute-based encryption and zero-knowledge proof techniques so that the updates initially come from manufacturer can be shared among the autonomous vehicles as a peer-to-peer manner, at the same time preserving the privacy as well as being targeted distribution. In the same connection, Puggioni et al. [52] proposes a IoT updates delivery approach by introducing a cryptocurrency and a rewarding scheme, integrated with zero-knowledge proof to address the scalability of typical update approaches. Meanwhile, a firmware update platform is proposed in [53] for IoT devices via hardware platform module, batch verification, digital signature, and anti-malware checking.
@@
$\times$
$\times$
$\times$
$\times$
$\times$
$\times$
$\times$
@@
$\times$
$\times$
$\times$
$\times$
$\times$
$\times$
@@ VII Conclusion and Future Works
However, our presented architecture will be flexible enough in terms of exploring and applying other cryptographic schemes as well as blockchain consensus mechanisms which will be solely depending on the different security requirements and application scenarios. Perhaps, in future work, the research direction can be as follows: (i) investigate how to integrate offchain concepts in untrusted scenarios with zero knowledge proof, (ii) look for mitigating the adversarial attacks against edge computing nodes, (iii) explore to incorporate Self-Sovereign Identity and data capsule techniques, (iv) design and apply a signcryption scheme having decentralization registration capabilities, and/or lesser storage & computational requirements, and (v) develop an anonymous reward payment system.
@@ Acknowledgments
The authors would like to thank to the editor and reviewers for thoughtful comments and constructive suggestions towards improving the quality of this work.

### v2

@@ Abstract
Smart devices are considered as an integral part of Internet of Things (IoT), have an aim to make a dynamic network to exchange information, collect data, analysis, and make optimal decisions in an autonomous way to achieve more efficient, automatic, and economical services. Message dissemination among these smart devices allows adding new features, sending updated instructions, alerts or safety messages, informing the pricing information or billing amount, incentives, and installing security patches. On one hand, such message disseminations are directly beneficial to the all parties involved in the IoT system. On the other hand, due to remote procedure, smart devices, vendors, and other involved authorities might have to meet a number of security, privacy, and performance related concerns while disseminating messages among targeted devices. To this end, in this paper, we design STarEdgeChain, a security and privacy aware targeted message dissemination in IoT to show how blockchain along with advanced cryptographic techniques are devoted to address such concerns. In fact, the STarEdgeChain employs a permissioned blockchain assisted edge computing in order to expedite a single signcrypted message dissemination among targeted groups of devices, at the same time avoiding the dependency of utilizing multiple unicasting approaches. Finally, we develop a software prototype of STarEdgeChain and show it’s practicability for smart devices.
@@ I Introduction
The Internet of Things (IoT) [1, 2, 3], has been rapidly growing more than ever before. From consumer electronic devices to industry, the IoT is making it possible to connect a variety of things and people to the Internet, and each other to bring enhancements to our quality of lives, surroundings, and system performances. Such things, also called smart devices (SDs) connect to the Internet and each other to exchange information, collect data, analyze, and make better decisions in an autonomous way without human operation. These abilities make SDs very popular in various application areas in the home, smart power grid, building, city, industry, healthcare, and many more. In specific, personal and home devices connected with IoT [4, 5] have been receiving much attention in the consumer electronics landscape since it is expected that the IoT is one of the key enablers to bring simplicity and ubiquitous connectivity to the consumers.
@@
Paper Organization: The remainder of this paper can be summarized as follows. Section II describes the problem statement including system settings, adversary model, and security requirements. The following section, the preliminary background of our construction is considered. Next, section IV presents the detailed construction of the proposed STarEdgeChain. After that, section V summarizes the security and performance analysis. Section VI discusses the related works. Finally, section VII concludes the presented work. The list of notations and their definitions used in this paper is described in Table I.
@@
$\lambda$ Security parameter
$p$ Prime number
$\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{3}$ Three cyclic groups
$e$ Bilinear map
$g_{1},g_{2}$ Generator of $\mathbb{G}_{1}$, generator of $\mathbb{G}_{2}$
$\mathbb{Z}_{p}$ Integers modulo $p$
$pk$ Public key
$sk$ Asymmetric secret key
$mk$ Master key
$key_{sym}$ Symmetric secret key
$key_{sign}$ Signing key
$key_{ver}$ Verification key
$msg$ Plaintext message
$\mathcal{S}$ The set of attributes
$\mathcal{T}$ Access tree structure
$\mathcal{A}$ An adversary
$\mathcal{ST}$ Signed Ciphertext
$\mathcal{CT}_{msg}$ Ciphertext of message
$\{0,1\}^{*}$ Set of binary strings
$\mathcal{Y}$ Set of leaf nodes in access tree
$\mathcal{H}_{1},\mathcal{H}_{2}$ Hash functions
$\pi$ Signature
[…]
$\oplus$ XOR operation
@@ III-D Permissioned Blockchain and Consensus
Blockchain [28, 29, 30, 31, 32, 33], a tamper-proof and consensus based distributed ledger technique, consists of a chain of blocks. The security and privacy of blockchain is ensured by cryptographic techniques, consensus, and incentive mechanisms. Blockchain offers data auditability by employing authenticated blocks attached in the growing chain. In the chain, each block keeps records of financial transactions or non-financial data, and a number of blocks form a chain by including the cryptographic hash of the previous block.
@@ IV-D Message Dissemination
A discussion on the proposed permissioned blockchain architecture is as follows in terms of the following two point of views.
@@ V-B Implementation and Performance Evaluation
Implementation Environment: The service providers, PoA validators, and edge computing nodes commonly have high computational devices relatively to IoT smart devices. Thus, as presented in Fig. 3, we have taken two laptop computer, representing as two host machines (Acer Predator Triton 500 and Lenovo IdeaPad 320S) as service providers, PoA validators, and edge computing nodes, whereas, one Raspberry Pi device (RPi 3 Model B+) as IoT smart device. In the host machines, we have configured Oracle virtual box (https://www.virtualbox.org/) and installed Ubuntu 18.04 with one core, RAM 4GB as virtual machines. In fact, we have considered the aforementioned virtual machines and one smart device as node 1, 2, and 3, respectively.
Besides, we have used Python 3.6 version for developing the software prototype. After that, we have used the bridge network feature from Oracle virtual box and made a local network among the host machines, virtual machines, and the RPi device for the purpose of on-chain implementation. We have also developed a graphical user interface (GUI) for convenience operation. Specifically, in Fig. 3, the node 1 represents a service provider making transactions and doing PoA self-validation of transactions to add in the chain. On the other hand, node 2 shows an edge computing node connected with a blockchain network and supported to perform targeted message dissemination to the RPi device considered as IoT smart device in node 3.
The codes and implementation instructions are publicly available to facilitate the reproducibility of this work at Github [^1] https://github.com/mbaqer/Blockchain-IoT. The Table II presents a summery of the computing nodes, assigned IP addresses & ports, used libraries, and other configurations. In particular, we have utilized Flask and Tkinter to develop a web based application and GUI, respectively, whereas the Charm-Crypto and hashlib are cryptography libraries, which are discussed in details in the next paragraph.
@@
Node 1 Intel Core i7-10th generation, RAM 32GB, and Microsoft Windows 10 operating system)
Node 2 Intel Core i5-8th generation, RAM 8GB, and Microsoft Windows 10 operating system)
Node 3 ARM Cortex-A53 (ARMv8) 1.4GHz quad-core processor, 1GB RAM, and Raspberry Pi OS)
Libraries Flask, Tkinter, Charm-Crypto, hashlib
IP and Port addresses Node 1 VM1: 192.168.100.10:5000,
@@
Overall, the computational costs introduced by the cryptographic processing in our developed permissioned blockchain are very modest for the lightweight IoT smart devices. Usually, the lightweight IoT smart device nodes have very limited storage to keep the entire blockchain and update it over time. Therefore, the interesting point is that in our permissioned blockchain, the lightweight nodes do not require to store the entire blockchain, rather the nodes will be relied on edge computing nodes. This shows the advantages of employing this work into wearable devices (smartwatches and fitness trackers), smart home devices (smart speakers, thermostats, security cameras, lights, and appliances), and healthcare IoT gadgets (blood pressure monitors, glucose meters, and ECG devices like personal health monitoring devices), just to name a few.
@@ VI Related Works
Typical Message Dissemination: In recent years, many solutions on typical message dissemination approaches have been introduced. For instance, the work in [38] focuses on secure and privacy-aware emergency vehicular message dissemination named as BTMPP. This solution utilizes bloom filter enabled private set intersection techniques to offer strong conditional privacy. Also, this proposed approach supports precise trust management simultaneously. Moreover, the authors in [39] proposes a privacy preserving trust management approach for message dissemination in vehicular space-air-ground integrated networks, which has an aim to ensure strong applicability and robustness by employing 0-1 encoding approach. A novel low-latency cooperative security method presents in [40] for platoon-assisted vehicular network to ensure timely and secure vehicular data dissemination. Here, a secret key agreement scheme is employed, in which the secret keys are generated based on the correlation of vehicles radio channels. The work in [41] introduces an end-to-end fast and secure firmware update framework IoT device like embedded systems. This work mainly employed a hardware-embedded cryptographic primitive referred to as a physical unclonable function (PUF) technique. Another work presented in [42] proposes a remote attestation-based firmware update for IoT. Kornaros et al. [43] present a secure firmware update for connected vehicles by including timestamp, version of update as well signed metadata of update.
@@
Integration of Blockchain: In terms of blockchain integration in message dissemination, existing works mainly employ blockchain along with different cryptographic schemes. In this context, cryptographic techniques are utilized to guarantee the security and privacy related requirements to the involved entities and message contents. For example, the work in [49] constructs a public key infrastructure (PKI) based private blockchain to offer privacy-aware message dissemination in vehicular networks. Another work in [50], the authors develop a semi-decentralized attribute-based blockchain to distribute the traffic signal messages and assist traffic control among the vehicles with high availability and privacy preserving ways in IoV. Here, the authors utilize blockchain along with a ciphertext-policy ABE scheme. Besides, the authors in [51] and [52] presents on how typical blockchain structure is able to handle data dissemination in Industrial IoT and Internet of Drones, respectively. Moreover, Baza et al. [53] leverage attribute-based encryption and zero-knowledge proof techniques so that the updates initially come from manufacturer can be shared among the autonomous vehicles as a peer-to-peer manner, at the same time preserving the privacy as well as being targeted distribution. In the same connection, Puggioni et al. [54] propose a IoT updates delivery approach by introducing a cryptocurrency and a rewarding scheme, integrated with zero-knowledge proof to address the scalability of typical update approaches. Meanwhile, a firmware update platform is proposed in [55] for IoT devices via hardware platform module, batch verification, digital signature, and anti-malware checking.
@@ VII Conclusion and Future Works
However, our presented architecture will be flexible enough in terms of exploring and applying other cryptographic schemes as well as blockchain consensus mechanisms which will be solely depending on the different security requirements and application scenarios. Perhaps, in future work, the research direction can be as follows: (i) investigating how to integrate offchain concepts in untrusted scenarios with zero knowledge proof, (ii) looking for mitigating the adversarial attacks against edge computing nodes, (iii) exploring how to incorporate Self-Sovereign Identity and data capsule techniques, (iv) designing and applying a signcryption scheme having decentralization registration capabilities, and/or lesser storage & computational requirements, and (v) developing an anonymous reward payment system.
@@ Acknowledgments
The authors would like to thank to the editor and reviewers for thoughtful comments and constructive suggestions toward improving the quality of this work.